WAVEGUIDE CHIRPED-PULSE FOURIER TRANSFORM MICROWAVE SPECTROSCOPY OF 2-ETHOXYETHANOL

\underline{MARIA A. PHILLIPS} and STEVEN T. SHIPMAN, Division of Natural Sciences, New College of Florida, Sarasota, FL 34243.

The pure rotational spectrum of 2-ethoxyethanol was recorded from 8.7 to 26.5 GHz at 250 K with a waveguide chirpedpulse Fourier transform microwave spectrometer. The full spectrum contains contributions from multiple vibrational states. Preliminary assignments have been made with a combination of ab initio calculations and an automated spectral fitting program that accelerates the initial fitting process.